OpenStudy (anonymous):

PLEASE HELP ME! Solve for x in this equation. 1/2= 14/6x-2 - 6/3x-1 pleaseeeee help me.

OpenStudy (anonymous):

\[\frac{1}{2}=\frac{14}{6x-2}-\frac{6}{3x-1}\]?

OpenStudy (anonymous):

yes Nancy

OpenStudy (anonymous):

is that 1/2=14/(6x-2)-6/(3x-1) if so, then: 1=14/(3x-1)-12/(3x-1) (multiplied both sides by 2) 1=2/(3x-1) 3x-1=2 x=1

OpenStudy (anonymous):

OH MY GOD thank you sooo much Stroodle!

OpenStudy (anonymous):

I think u have answer

OpenStudy (anonymous):

@twilight: award the person with a medal by clicking the good answer tab on the right. (s)he has spent time to answer your question!

OpenStudy (anonymous):

yep x=1